Consumption under borrowing constraint

(a) With borrowing constraint, household can borrow until 200 in period 1. Under y1 =100, y2= 200, r= 0.2,

How much is maximum possible consumption of period 1?
How much is maximum possible saving of period 1? (under maximum possible of c1)
How much is maximum possible consumption of period 2? (under maximum possible of c1)

(b) With full borrowing constraint, household cannot borrow any money in period 1 and cannot consume as well. (C1 = 0)

Under y1 = 100, y2= 200, r= 0.2, How much is maximum possible consumption of period 2? Write budget constraint of period 2.
Answer #1

A) maximum possible Consumption in period 1,

= Y1 + maximum poseposs borrowing

= 100+200 = 300

Then maximum possible saving, = 0

Since with maximum possible C1 & borrowing upto 200, no saving is possible

So maximum C2 = 0

Bcoz all period 2 income is borrowed in period 1

B)

Maximum possible Consumption of period 2

= Y2 + (1+r) Y1

= 200+(1.2)100

= 200+120

= 320

Period 2 BC,

C2 = Y2+(1+r)Y1 = 320